OpenStudy (anonymous):

how would you subtract 3pi/4 from 2pi

zepdrix (zepdrix):

Since the first term is a fraction, we need to write the second term as a fraction also. You would need to get a common denominator.\[\large \frac{2\pi}{1}-\frac{3\pi}{4}\]See how I wrote the first term as a fraction? We'll multiply the first term by \(\dfrac{4}{4}\) to get a common denominator.

zepdrix (zepdrix):

\[\large \color{royalblue}{\frac{4}{4}}\cdot \frac{2\pi}{1}-\frac{3\pi}{4}\]

zepdrix (zepdrix):

\[\large\frac{8\pi}{4}-\frac{3\pi}{4}\]Then just subtract the tops! :)

OpenStudy (anonymous):

Thank You!